# Correcting Negative Bias in Large Language Models through Negative Attention Score Alignment

### v1

@@ 3.2.2 Empirical Studies
Model LLaMA3-8B Mistral-7B Gemma-7B Qwen2-7B
@@
Table 1: Correlations of NAS and the negative confidence ($-1$ to $1$).
To demonstrate that the previously defined NAS is an effective indicator of negative bias, we measure the correlation between NAS and the negative confidence observed in model responses. For this measurement, we compile a test set consisting of a total of 1,500 samples, drawing 500 samples each from StrategyQA, rephrased GSM8K, and AR-LSAT. Table 1 shows the Pearson correlation and Spearman’s rank correlation coefficient measured between NAS and negative response confidence for each model. In all cases, we confirm that NAS indeed has a positive correlation with negative confidence. This validates that NAS is an effective indicator of a model’s negative bias.
@@
Additionally, we conduct a study on negative heads defined through the Negative Attention Score (NAS). For the QA dataset composed of three different domains that we previously created, we explore negative attention heads in three cases based on each domain subset. We designate as negative attention heads for a domain subset those heads among the top 200 heads with the highest single head NAS per sample that appear commonly in over 90% of the subset samples. Keeping the model type constant, we extract such negative heads from each subset and then measure the extent of their overlap. The overlapping rates are displayed in Table 2, and surprisingly, we find that negative heads extracted through different domain subsets significantly overlap. This indicates the existence of query-agnostic negative attention heads that represent a common underlying cause of the model’s negative bias.
LLaMA3-8B Mistral-7B Gemma-7B Qwen2-7B
@@
Table 2: Overlapping proportion of top-200 negative heads for various datasets (00 to $1$).
4 NASA: Negative Attention Score Alignment
@@ 4.1 Probing Set Construction
Motivated by previous observations, we aim to quantify the negative bias of LLMs from the perspective of attention heads. To observe the negative bias in attention heads, we start by constructing a negative attention head probing set. To construct the probing set, we first select samples that can be answered by the model’s parametric knowledge then convert them into binary decision making format using Prompt I. Note that in the probing set, the label is uniformly set to “Yes" to effectively reveal the model’s negative bias.
We suppose that the probing samples must meet the condition: the question should be answerable using the model’s parametric knowledge. To determine whether the question in the sample meets the condition, we construct the prompt as follows:
This prompt effectively selects samples that belong to the model’s parametric knowledge while minimizing the influence of the model’s negative bias. We assume that samples meet the condition if the model returns the correct answers for the prompt, and we use them as the probing set. We utilize GPT-4 to verify whether the prediction matches to label. A detailed explanation of the parametric sample selection can be found in Appendix D.
In Section 3, we observe that the negative bias attention heads exhibit prompt-agnostic properties. Based on this observation, we employ another multi-step reasoning dataset, HotpotQA (Yang et al., 2018), to construct the probing set. We presume that HotpotQA, being less challenging compared to recent datasets, has a higher proportion of samples that can be answered using the parametric knowledge of LLMs. For the samples selected for probing, we construct the probing set by converting them into the form of binary decision prompt (See Section 3.1.1).
@@ 4.2 Negative Attention Head Probing
We probe negatively biased attention heads using the constructed probing set. First, we measure the NAS of each attention head when the prompt from each probing sample is input into the model. We then identify the top $k^{\prime}$ attention heads with the highest NAS for each probing sample. Among these extracted heads, we select those that are consistently extracted in over 90% of the probing set samples. Finally, the top $k$ attention heads, sorted by their single head NAS values, are targets for further addressing. In this paper, we set $k^{\prime}$ to 100.
At this point, we further categorize the probing set into cases where the model correctly answers (i.e., the true positive set) and cases where it does not (i.e., the false negative set). During the fine-tuning phase of the model (See Section 4.3), we use the false negative set as the training set. This approach is aimed at inducing the model to generate positive answers for samples that it incorrectly fails to answer due to a negative bias in parametric knowledge. Meanwhile, the true positive set is used to determine the criteria for early halting of the fine-tuning process. In other words, during the training process aimed at reducing the model NAS, if the NAS on the validation set (i.e., false negative subset) falls below the minimum model NAS of true positive set, training is halted to minimize the bias towards positives.
@@ 4.3 Head-wise Incremental Tuning
To address the identified attention heads as specified in Section 4.2, we propose an additional supervised fine-tuning method. Our fine-tuning method is parameter efficient as it focuses only on fine-tuning the query and key projection weights of the selected negative bias attention heads.
4.3.1 Dataset
We utilize the data from the attention head probing set for fine-tuning (i.e., subset of rephrased HotpotQA). Because the model already possesses parametric knowledge for answering questions in the probing set, we construct the fine-tuning data to encourage the model to generate positive answers in response to Prompt I. As the queries in the probing set contain parametric knowledge, we do not provide contexts containing related facts to the model.
Dataset Model Accuracy Precision Recall F1 NAS
StrategyQA LLaMA3-8B / + NASA 0.871 / 0.877 0.919 / 0.864 0.795 / 0.875 0.852 / 0.869 160.6 / 55.0
Mistral-7B / + NASA 0.842 / 0.830 0.882 / 0.802 0.772 / 0.861 0.823 / 0.830 114.1 / 34.3
Gemma-7B / + NASA 0.778 / 0.771 0.923 / 0.907 0.676 / 0.690 0.777 / 0.784 107.8 / 71.8
Qwen2-7B / + NASA 0.829 / 0.844 0.921 / 0.907 0.694 / 0.742 0.792 / 0.816 209.6 / 107.2
MuSiQue (Rephrased) LLaMA3-8B / + NASA 0.720 / 0.757 0.859 / 0.815 0.555 / 0.683 0.675 / 0.743 199.8 / 64.0
Mistral-7B / + NASA 0.649 / 0.707 0.848 / 0.802 0.416 / 0.594 0.558 / 0.682 131.2 / 110.7
Gemma-7B / + NASA 0.518 / 0.503 0.837 / 0.834 0.529 / 0.555 0.648 / 0.667 259.8 / 96.2
Qwen2-7B / + NASA 0.621 / 0.639 0.883 / 0.858 0.482 / 0.552 0.624 / 0.672 418.8 / 282.9
@@ 4.3.1 Dataset
GSM8K (Rephrased) LLaMA3-8B / + NASA 0.537 / 0.545 0.587 / 0.531 0.336 / 0.838 0.428 / 0.650 218.2 / 74.7
Mistral-7B / + NASA 0.501 / 0.512 0.518 / 0.509 0.163 / 0.782 0.247 / 0.617 115.5 / 77.0
Gemma-7B / + NASA 0.533 / 0.524 0.546 / 0.541 0.837 / 0.887 0.661 / 0.672 159.6 / 66.2
Qwen2-7B / + NASA 0.538 / 0.576 0.640 / 0.614 0.534 / 0.738 0.582 / 0.670 210.9 / 82.9
MATH (Rephrased) LLaMA3-8B / + NASA 0.563 / 0.549 0.587 / 0.534 0.530 / 0.874 0.557 / 0.663 183.3 / 60.8
Mistral-7B / + NASA 0.504 / 0.526 0.527 / 0.522 0.136 / 0.685 0.216 / 0.592 117.9 / 109.6
Gemma-7B / + NASA 0.546 / 0.543 0.556 / 0.553 0.810 / 0.871 0.659 / 0.677 185.0 / 64.0
Qwen2-7B / + NASA 0.564 / 0.581 0.661 / 0.617 0.537 / 0.715 0.593 / 0.662 224.0 / 103.1
AR-LSAT (Rephrased) LLaMA3-8B / + NASA 0.513 / 0.518 0.547 / 0.537 0.297 / 0.444 0.385 / 0.486 178.1 / 63.6
Mistral-7B / + NASA 0.511 / 0.530 0.547 / 0.553 0.280 / 0.447 0.370 / 0.494 108.7 / 80.02
Gemma-7B / + NASA 0.509 / 0.510 0.525 / 0.523 0.468 / 0.534 0.495 / 0.528 151.7 / 68.0
Qwen2-7B / + NASA 0.492 / 0.496 0.529 / 0.528 0.110 / 0.182 0.182 / 0.270 317.6 / 207.3
@@
4.3.2 Training Objective
Similar to a standard supervised fine-tuning objective, we target only the answer token (e.g., “Yes”) for training. Since our dataset contains only short answers, training the model to generate the special token that indicates the end of a sequence could introduce a bias towards short responses. To prevent this side effect, we only use the target answer token in our loss calculation.
4.3.3 Training Strategy
We propose to further fine-tune each attention head sequentially in order of decreasing NAS which is measured during the attention head probing stage. Specifically, we only train the query and key projection modules of the target attention head.
Since we train with data where the target answer is positive, there is a risk of inducing a positive bias. Furthermore, updating a single attention head might change the NAS of another attention head. To prevent this, we set aside a portion of the dataset as a validation set and apply early stopping and update cancellation schemes for each attention head tuning, and early halting scheme for the whole fine-tuning process.
Strategy of attention head tuning subprocess
we set early stopping criteria based on the model NAS and single head NAS. For the former criteria, we check for a decrease in the value each epoch and halt training if there is an increase. For the latter one, we verify that i) the value decreases each epoch and ii) remains above a specific threshold. Training is also stopped if either of the two conditions is not met.
After fine-tuning each attention head, we calculate the difference between the single NAS and the whole NAS before and after training the attention head. If one or both of these values have increased compared to before training, we revert the parameter update of that attention head. This is to exclude modules that reinforce the negative bias of other attention heads from the training target.
Strategy of the whole tuning process
In Section 4.2, we measure the NAS of each sample within the true positive probing set to establish the minimum value as the early halting threshold, $\tau$. We hypothesize that if the NAS value in a false negative sample is lower than in a true positive, it may induce a positive bias. After training each negative bias attention head, we measure the NAS across all attention heads in the validation set. If this value falls below $\tau$, training is halted to prevent falling into local optima.
@@ 5.1 Experimental Setup
We conduct experiments using LLaMA3, Mistral, Gemma, and Qwen2. We use those instruct models before further NASA fine-tuning as baselines. We use LLaMA-Factory (Zheng et al., 2024) [^1] https://github.com/hiyouga/LLaMA-Factory and HuggingFace Transformers (Wolf et al., 2020) [^2] https://github.com/huggingface/transformers for all the experiments. The models are evaluated using yes-no QA datasets, which are described in Section 3. For these datasets, we measure accuracy, precision, recall, F1, and the sum of NAS values across all attention heads in the attention head probing set. For early stopping, we halt training if two NAS values (refer to the last paragraph in Section 4.3) cease to decrease or if the NAS for the target attention head on the validation set drops below 0.5. The initial learning rate is set to 1e-6, the batch size to 32, and the maximum number of epochs to 30. We reference the Alpaca repository [^3] https://github.com/tatsu-lab/stanford_alpaca, setting the weight decay to 0 and the warmup ratio to 0.03.
@@ 5.2 Main Results
In this section, we examine whether the negative bias observed and analyzed in Section 3 has been effectively mitigated by NASA. Tables 3 and 4 present the performance of our method across various binary decision tasks. Generally, we observe an improvement in accuracy and a significant reduction in the gap between precision and recall. As mentioned in Section 3, high precision coupled with low recall indicates that the model is excessively cautious in providing positive responses. This, in turn, leads to a degradation in the trustworthiness of negative responses.
From this perspective, reducing the gap between precision and recall while maintaining accuracy at an equivalent or higher level is a way to mitigate negative bias, balancing the trustworthiness of the model’s responses, and ultimately improving the model. Additionally, based on the improvement in the F1 score, which is the harmonic mean of precision and recall, it is demonstrated that the balance between precision and recall has been achieved in a positive direction of improving the model capability for reasoning. Regarding the NAS metric, we observe that our method effectively reduces NAS to an appropriate level. In Appendix G, we demonstrate that the components of our method contribute effectively to consistently tackling NAS during the training process.
Meanwhile, for the mathematical reasoning task, recall surpasses precision. This indicates that this task is somewhat more sensitive compared to other tasks and suggests that additional regularization, considering the task’s sensitivity, could enhance the effectiveness of our method. Future research based on our work could delve deeper into this aspect.
Moreover, the slight decrease in precision might superficially appear as a performance decline. However, it can be interpreted as an enhancement of the model’s trustworthiness by adjusting the alignment between the model’s actual knowledge and its responses. This means that precision might slightly decrease because some samples that were previously incorrectly categorized as true negatives due to negative bias might shift to false positives. For instance, even if the model makes incorrect inferences or remains uncertain about a question, in a binary decision scenario, negative bias could lead to it being classified as a true negative. A detailed analysis of the shifts in negative responses due to NASA can be found in Appendix F.
@@ 6 Analysis
6.1 Preservation of General Reasoning Abilities

### v2

@@ 3.2.2 Empirical Studies
Table 1: Correlations of NAS and the negative confidence ($-1$ to $1$).
Model LLaMA Mistral Gemma Qwen
@@
NAS and Negative Confidence To demonstrate that the previously defined NAS is an effective indicator of negative bias, we measure the correlation between model NAS and the negative confidence observed in model responses. For this measurement, we construct a test set consisting of a total of 1,500 samples, drawing 500 samples each from StrategyQA, rephrased GSM8K, and rephrased AR-LSAT. Table 1 shows the Pearson correlation and Spearman’s rank correlation coefficient measured between model NAS and negative response confidence for each model. In all cases, we confirm that model NAS has a positive correlation with negative confidence. This validates that NAS is an effective indicator of a model’s negative bias.
@@
Table 2: Overlapping proportion of top-100 negative heads for various datasets (00 to $1$).
LLaMA Mistral Gemma Qwen
@@
Overlapping Negative Attention Heads We conduct a study on negative attention heads defined through NAS. Using our previously constructed test set, which spans three different domains, we investigate whether the negative attention heads extracted from each domain subset overlap with one another. For each sample $x_{i}$ in a domain subset $X$, we first define the tuple list $P^{i}$ consisting of the layer and head indices of the top-$k$ attention heads with the highest single head NAS:
[…]
Among these extracted heads, we select those that are consistently included in over 90% of the tuple lists, which we denote as $C_{X}$:
$C_{X}=\left\{(l,h)\ \middle|\right.$
$\left.(l,h)\text{ appears in at least }0.9\times|X|\text{ of }P_{X}^{i}\right\}.$ (3)
Finally, we extract the list of top $N$ attention heads $P_{X}$, sorted by their single head NAS values on $X$.
[…]
In this process, we set $k$ to 200 and $N$ to 100.
Keeping the model type constant, we extract such negative attention heads from each domain subset and then measure the extent of their overlap. The overlapping rates are displayed in Table 2, and surprisingly, we find that negative heads extracted through different domain subsets significantly overlap. This indicates the existence of query-agnostic negative attention heads that represent a common underlying cause of the model’s negative bias.
4 NASA: NAS Alignment Framework
Motivated by previous observations, we propose NAS Alignment (NASA), a framework designed to address attention heads that induce negative bias in a parameter-efficient manner. In NASA, we start by constructing a probing set to identify negative attention heads, followed by a selection process to determine which of these heads will be fine-tuned. Our framework fine-tunes the target attention heads in the order based on single head NAS.
@@ 4.1 Probing Set Construction
To construct the probing set, we use a short-answer QA dataset and select the parametric samples where the model can correctly answer the question. A detailed explanation of the parametric sample selection can be found in Appendix E. Next, we convert the selected parametric samples into binary decision making format using Prompt I in Section 3.1.1. Note that in the probing set, the label is uniformly set to “Yes”. The probing set is designed to have two properties: i) the model possesses knowledge about the question, and ii) the model should respond “Yes” to the converted question. We assume that attention heads with a high single head NAS for the converted samples that meet the two conditions are strongly associated with negative bias and should be addressed.
In Section 3.2.2, we observe that the negative bias attention heads exhibit query-agnostic properties. Based on this observation, we employ another multi-step reasoning dataset, HotpotQA (Yang et al., 2018), to construct the probing set. We presume that HotpotQA, being less challenging compared to recent datasets, has a higher proportion of samples that can be answered using the parametric knowledge of LLMs.
@@ 4.2 Negative Attention Head Probing
Using the negative attention head detection process described in Section 3.2.2, we extract the attention heads to be fine-tuned by utilizing the probing set $X$. In this process, we set $k$ to 100 and $N$ to 30.
At this point, we further categorize the probing set $X$ into cases where the model correctly answers $TP(X)$ (i.e., the true positive set) and cases where it does not $FN(X)$ (i.e., the false negative set). During the fine-tuning phase of the model (See Section 4.3), we use the false negative set as the fine-tuning set $D$. This approach is aimed at inducing the model to generate positive answers for samples that it fails to answer due to a negative bias. Meanwhile, the true positive set is used to determine the threshold $\tau$ for early halting of the fine-tuning process.
$\tau:=\min_{x\in TP(X)}\text{NAS}(\{x\},L,H)$ (5)
In other words, during the training process aimed at reducing the model NAS, if the model NAS on the validation set (i.e., subset of $FN(X)$) falls below the minimum model NAS of true positive set, training is halted to minimize the bias towards positives.
@@ 4.3 Head-wise Incremental Tuning
Training setup
As mentioned in Section 4.2, we utilize the $FN$s in the probing set for fine-tuning. Because the model fails to generate positive answers for the samples although the model possesses parametric knowledge, it is expected that fine-tuning with these samples can effectively address the negative bias. We do not provide contexts containing related facts to the model because the model already has the corresponding knowledge.
@@
StrategyQA LLaMA / + NASA 0.871 / 0.877 0.919 / 0.864 0.795 / 0.875 0.852 / 0.869 160.6 / 55.0
Mistral / + NASA 0.842 / 0.830 0.882 / 0.802 0.772 / 0.861 0.823 / 0.830 114.1 / 34.3
Gemma / + NASA 0.778 / 0.771 0.923 / 0.907 0.676 / 0.690 0.777 / 0.784 107.8 / 71.8
Qwen / + NASA 0.829 / 0.844 0.921 / 0.907 0.694 / 0.742 0.792 / 0.816 209.6 / 107.2
MuSiQue (Rephrased) LLaMA / + NASA 0.720 / 0.757 0.859 / 0.815 0.555 / 0.683 0.675 / 0.743 199.8 / 64.0
Mistral / + NASA 0.649 / 0.707 0.848 / 0.802 0.416 / 0.594 0.558 / 0.682 131.2 / 110.7
Gemma / + NASA 0.518 / 0.503 0.837 / 0.834 0.529 / 0.555 0.648 / 0.667 259.8 / 96.2
Qwen / + NASA 0.621 / 0.639 0.883 / 0.858 0.482 / 0.552 0.624 / 0.672 418.8 / 282.9
@@
Dataset Model Accuracy Precision Recall F1 NAS
GSM8K (Rephrased) LLaMA / + NASA 0.537 / 0.545 0.587 / 0.531 0.336 / 0.838 0.428 / 0.650 218.2 / 74.7
Mistral / + NASA 0.501 / 0.512 0.518 / 0.509 0.163 / 0.782 0.247 / 0.617 115.5 / 77.0
Gemma / + NASA 0.533 / 0.524 0.546 / 0.541 0.837 / 0.887 0.661 / 0.672 159.6 / 66.2
Qwen / + NASA 0.538 / 0.576 0.640 / 0.614 0.534 / 0.738 0.582 / 0.670 210.9 / 82.9
MATH (Rephrased) LLaMA / + NASA 0.563 / 0.549 0.587 / 0.534 0.530 / 0.874 0.557 / 0.663 183.3 / 60.8
Mistral / + NASA 0.504 / 0.526 0.527 / 0.522 0.136 / 0.685 0.216 / 0.592 117.9 / 109.6
Gemma / + NASA 0.546 / 0.543 0.556 / 0.553 0.810 / 0.871 0.659 / 0.677 185.0 / 64.0
Qwen / + NASA 0.564 / 0.581 0.661 / 0.617 0.537 / 0.715 0.593 / 0.662 224.0 / 103.1
AR-LSAT (Rephrased) LLaMA / + NASA 0.513 / 0.518 0.547 / 0.537 0.297 / 0.444 0.385 / 0.486 178.1 / 63.6
Mistral / + NASA 0.511 / 0.530 0.547 / 0.553 0.280 / 0.447 0.370 / 0.494 108.7 / 80.02
Gemma / + NASA 0.509 / 0.510 0.525 / 0.523 0.468 / 0.534 0.495 / 0.528 151.7 / 68.0
Qwen / + NASA 0.492 / 0.496 0.529 / 0.528 0.110 / 0.182 0.182 / 0.270 317.6 / 207.3
Similar to a standard supervised fine-tuning objective, we target only the answer token (e.g., “Yes”) for training. Since the fine-tuning set contains only short answers, training the model to generate the end of a sequence token could introduce a bias towards short responses. To prevent this side effect, we exclude the end of a sequence token in a loss calculation.
Training strategy
We fine-tune the query and key projection modules of each attention head sequentially in order of decreasing single head NAS which is measured during the probing stage.
Since we fine-tune with data where the target answer is positive, there is a risk of inducing a positive bias. Furthermore, updating a single attention head might change the NAS of another attention head. To prevent these, we set aside a portion of the dataset as a validation set and apply early stopping and update cancellation schemes for each attention head tuning and early halting scheme for the whole fine-tuning process.
For early stopping, we set criteria based on the model NAS and single head NAS. For the former criteria, we check for a decrease in the value after each epoch. For the latter one, we verify that i) the value decreases and ii) remains above a specific threshold $\rho$ after each epoch. Fine-tuning of the current attention head is stopped if either of the three conditions is not met.
For the update cancellation, after fine-tuning each attention head, we calculate the difference between the single head NAS and the model NAS before and after fine-tuning. If any of these values have increased compared to before training, we revert the parameter update of that attention head to exclude attention heads that reinforce the negative bias from the training target.
For the early halting, we utilize the model NAS of each sample within the $TP(X)$, $\tau$, as mentioned in Section 4.2. This threshold setting is based on the rationale that if the model NAS in a $FN(X)$ is lower than in a $TP(X)$, it may induce a positive bias. After training each negative attention head, we measure the model NAS on the validation set. If this value falls below $\tau$, training is halted to prevent falling into local optima. Algorithm for the entire process can be found in Appendix F.
@@ 5.1 Experimental Setup
We use LLaMA, Mistral, Gemma, and Qwen for our experiments. We use those supervised fine-tuned models before NASA-tuning as baselines. We use LLaMA-Factory (Zheng et al., 2024a) [^1] https://github.com/hiyouga/LLaMA-Factory and HuggingFace Transformers (Wolf et al., 2020) [^2] https://github.com/huggingface/transformers for all the experiments. The models are evaluated using binary decision datasets, which are described in Section 3, measuring accuracy, precision, recall, F1, and the model NAS. We set $\rho$ to 0.5, the initial learning rate to 1e--6, the batch size to 32, and the maximum number of epochs to 30. We reference the Alpaca repository [^3] https://github.com/tatsu-lab/stanford_alpaca, setting the weight decay to 0 and the warmup ratio to 0.03.
@@ 5.2 Main Results
Tables 3 and 4 show the performance of our method across various binary decision tasks. Generally, we observe an improvement in accuracy and a significant reduction in the precision-recall gap. As noted in Section 3, high precision coupled with low recall indicates that the model is excessively cautious in providing positive responses, resulting in a degradation in the trustworthiness of negative responses. From this perspective, the negative bias can be mitigated by reducing the gap between precision and recall while maintaining accuracy, which balances the trustworthiness of the model’s responses and ultimately improves the model.
Additionally, based on the improvement in the F1 score, which is the harmonic mean of precision and recall, we demonstrate that the balance between precision and recall has been achieved in the direction of improving the model’s capability for reasoning. We also observe that our method effectively reduces the model NAS. In Appendix K, we demonstrate that the components of our method contribute effectively to tackling negative bias.
Meanwhile, the slight decrease in precision might superficially appear as a performance decline. However, it can be interpreted as enhancing the model’s trustworthiness by adjusting the alignment between the model’s actual knowledge and its responses. This means that precision might slightly decrease because some samples incorrectly categorized as true negatives due to negative bias might shift to false positives. Specifically, even if the model makes incorrect inferences or remains uncertain about a question, in a binary decision scenario, negative bias could lead to it being classified as a true negative. An analysis of the shifts in negative responses due to NASA can be found in Appendix H.
@@ 6 Analysis
We analyze the impact of NASA on the general reasoning capability of LLMs beyond the binary decision task which is the primary focus of this work. After that, we examine the effect of NASA on model calibration. Additionally, regarding performance improvement and negative bias, we investigate the relationship between few-shot examples and NASA in few-shot prompting scenarios. We present an instruction following analysis and an ablation study in Appendices I, J and K.
6.1 General Reasoning Abilities
We evaluate the performance of LLMs fine-tuned with NASA on general QA reasoning tasks. We utilize general reasoning benchmarks such as MuSiQue, GSM8K, and AR-LSAT without the transformation. For short-answer QA, we first verify whether each model-generated prediction includes an answer in these benchmarks and then utilize GPT-4 to double-check whether the prediction is semantically correct. To generate predictions from the model, we utilize Prompt II in Section 4.1. The content of the GPT-4 prompt for prediction verification can be found in Table 15.
As shown in Table 5, models fine-tuned with NASA show enhanced or maintained general reasoning abilities while effectively tackling negative bias. Since negative bias is a task-specific factor contributing to hallucination in binary decision tasks, its influence on general reasoning QA, like short-answer QA, is relatively low compared to other factors contributing to hallucination such as parametric knowledge. Therefore, improving reasoning ability in binary decision tasks does not necessarily enhance reasoning capabilities in general reasoning QA. This means that addressing bias in binary decision tasks while maintaining performance in general reasoning QA can be considered an advancement in the model’s overall reasoning capability, not overfitted to the binary decision task.
6.2 Enhanced Model Calibration
As Fig 2 illustrates, the negative bias of LLMs is also associated with the model’s confidence in its predictions. This issue regarding the gap between prediction confidence and actual accuracy is related to research on calibration, which aims to enhance the trustworthiness of LLMs (Kadavath et al., 2022). Table 6 shows that the expected calibration error (He et al., 2022) is reduced when NASA is applied to Mistral. This indicates that NASA effectively mitigates model bias, better aligning prediction confidence and accuracy.